Can someone help me solve this?
Solve the system by the substitution method.
4x+2y=-8
2x+y=-4

Answer by jim_thompson5910(35256)

Solve the second equation for y

2x+y=-4

y = -4-2x

Then use this back in the first equation

4x+2y=-8

4x+2(-4 - 2x)=-8 ... replace y with -4-2x (we've just shown that y = -4-2x)

4x-8-4x=-8

0x-8 = -8

0-8 = -8

-8 = -8

Since this last equation is ALWAYS true (regardless of what x is), this means that there are an infinite number of solutions

So this system is consistent and dependent.